What is the heat required to boil 743 grams of water in a kettle at 100c


1
Expert's answer
2021-02-03T07:15:04-0500

Solution:Given:

"m=743g"

"c=2.02\;(\frac{J}{g\;C})"

"\Delta T=100^oC"


"Q=mc\Delta T=(743g)(2.02 \;(\frac{J}{g\;C}))(100^oC)=150086J=150.086kJ"